OpenStudy (anonymous):

find three consecutive integers whose sum is 363

OpenStudy (anonymous):

120+121+122=363

OpenStudy (anonymous):

Please give medal

OpenStudy (anonymous):

let the numbers be x.x+1,x+2 x+x+1+x+2=363 calculate x and then x+1 and x+2

OpenStudy (anonymous):

that equals 366 NOT 363 @wsammixboo

sammixboo (sammixboo):

Oh sorry xD

OpenStudy (anonymous):

no prob that was my first guess too, then i changed it xD

sammixboo (sammixboo):

119+121+123=363 Hahah!

OpenStudy (anonymous):

3x+3=363 3x=363-3=360 x=360/3=120 x+1=? x+2=?

OpenStudy (anonymous):

@surjithayer I understand your steps. It never occured to me to use this formula :D

OpenStudy (anonymous):

take any no.say 20, next no.=20+1 and next=20+2 similarly take any no. say x next no.=x+1 and next =x+2